# Adaptive Data-Knowledge Alignment in Genetic Perturbation Prediction

### v1

@@ 4.2 Knowledge Refinement of Gene Regulatory Networks (Q2)
We used the accuracy of interactions to test if ALIGNED’s knowledge refinement can re-discover underlying regulations from synthetic data generated from OmniPath GRN (Türei et al., 2016). For topology, we evaluated the method’s ability to produce well-structured GRNs, in terms of: 1) network modularity, for clustering quality of functional modules (Alon, 2007) and 2) degree assortativity, for regulatory hub structures (Segal et al., 2003). In addition, we examined the method’s ability to re-discover biologically meaningful interactions by cross-referencing external pathway databases. We compared overlaps with refined pathways using a gene set recovery algorithm (Huang et al., 2018) to obtain pathway enrichment scores.
@@
Our method was compared with a baseline using non-sparse (Frobenius norm) regularization. Existing approaches, such as GRN inference, treat the knowledge base as static instead of performing incremental refinement, and therefore are not suitable for the comparison.
@@
Observation 3. In Figure 5c, there are no significant differences of enrichment scores between the original and refined GRNs in most pathways. This indicates that ALIGEND can re-discover biologically meaningful knowledge annotated in cross-reference databases.
@@ 4.3 Perturbation Prediction on Bacterial Genome (Q3)
Setting, Dataset, and Knowledge Base. We evaluate our method on the Escherichia coli (E. coli) K-12 MG1655 strain using a combined dataset that includes 70 knockout perturbations by Lamoureux et al. (2023) (comprising 4 triple, 7 double, and 59 single perturbations, totaling 433 samples); and 7 data series with 16 single overexpression perturbations (73 samples) from the NCBI sequence read archive (Sayers et al., 2025). The knowledge base is constructed from the EcoCyc GRN (Moore et al., 2024), covering 315 regulator genes and 3,004 regulated genes. To evaluate generalization on unseen instances, we split the test set of unseen perturbations including 4 single overexpressions, 5 double knockouts, and 2 triple knockouts.
Similar to Section 4.1, we conducted ablation studies comparing ALIGNED against baseline models trained only on labelled data. We tracked performance through two complete iterations of ALIGNED to assess the cumulative effect of each framework component.
@@
Table 1: Performance of ALIGNED on E.coli genome.
Model ABL Stage Data Cons. Knowledge Cons. Balanced Cons.
MLP Baseline $0.3312_{\pm 0.0004}$ $0.3371_{\pm 0.0009}$ $0.3341_{\pm 0.0006}$
GNN Baseline $0.3773_{\pm 0.0026}$ $0.3605_{\pm 0.0024}$ $0.3689_{\pm 0.0012}$
ALIGNED
Align 1 $0.3872_{\pm 0.0009}$ $0.3708_{\pm 0.0006}$ $0.3784_{\pm 0.0002}$
Refine 1 $0.3872_{\pm 0.0009}$ $0.3836_{\pm 0.0067}$ $0.3851_{\pm 0.0033}$
Algin 2 $0.3812_{\pm 0.0029}$ $0.4025_{\pm 0.0020}$ $0.39404_{\pm 0.0018}$
Refine 2 $0.3812_{\pm 0.0029}$ $0.4045_{\pm 0.0038}$ $0.3912_{\pm 0.0013}$
ALIGNED
Algin 1 $0.3876_{\pm 0.0060}$ $0.3714_{\pm 0.0269}$ $0.3800_{\pm 0.0159}$
Refine 1 $0.3876_{\pm 0.0060}$ $0.4520_{\pm 0.0070}$ $0.4130_{\pm 0.0059}$
Align 2 $0.3878_{\pm 0.0064}$ $0.4668_{\pm 0.0235}$ $0.4124_{\pm 0.0042}$
Refine 2 $\mathbf{0.3878}_{\pm 0.0064}$ $\mathbf{0.5348}_{\pm 0.0069}$ $\mathbf{0.4288}_{\pm 0.0063}$
@@ 5 Related Work
Perturbation Response Prediction. Recent approaches fall into two categories: methods that utilize the compositional nature of genetic perturbation responses in learning latent representations (Lotfollahi et al., 2023; Cui et al., 2024; Hao et al., 2024), hybrid methods that leverage prior knowledge from biological networks (Roohani et al., 2024; Wenkel et al., 2025; Littman et al., 2025) or textual embeddings (Wang et al., 2024). In contrast, ALIGNED does not assume GRNs to be static and can systematically refine GRNs by adaptive learning from datasets and knowledge bases to leverage reliable information.
Neuro-Symbolic Learning. The Abductive Learning (ABL) framework (Zhou, 2019) integrates deep learning with symbolic constraints through consistency optimization. Extensions include $Meta_{abd}$ (Dai & Muggleton, 2021) for visual-symbolic reasoning, $ABL_{NC}$ for knowledge refinement (Huang et al., 2023), and $ABL_{refl}$ (Hu et al., 2025) for efficient neuro-symbolic integration using reinforcement learning mechanisms. Additionally, Cornelio et al. (2023) proposed a learnable trade-off mechanism between data and knowledge sources. We extend these approaches to biological systems where both experimental data and curated knowledge exhibit domain-specific noise and incompleteness.
@@ 6 Conclusion and Future Work
While we acknowledge the limitations in our regulatory network modelling, alternative methods (Covert et al., 2004; Stoll et al., 2017; Abou-Jaoudé et al., 2016) face significant scalability issues. Future work could explore differentiable models (Faure et al., 2023) and refine them with experimental data. Furthermore, ALIGNED can be extended to different biological tasks by leveraging other prior knowledge, such as protein-protein interaction networks (Rodriguez-Mier et al., 2025) and metabolic networks (Faure et al., 2023).
@@ Appendix A Usage of Large Language Models
We used Claude and ChatGPT mainly to polish the language after all intellectual content has been drafted, along with other language editing tools such as Grammarly.
@@ B.3 Discussion
In actual experiments in Section 4.1, we introduced an assumption that up/down regulation of a node can be decided by its in-degree of positive and negative interactions, in order to capture more detailed regulatory behaviours. This is achieved by using an integer variant of Equation 2.2. In this setting, the value of node 5 in Figure 7 will be -1 when $k=2$, i.e. $\delta_{\langle{\bm{R}}_{+}^{(2)},{\bm{R}}_{-}^{(2)}\rangle}({\bm{x}})=[0,-1,1,1,-1]^{\top}$, and the network behaviour will be more complicated in complex networks. However, such modelling is still not enough to describe the biological reality, and future work could further explore the other differentiable modelling approaches of genome-scale GRN under the ALIGNED framework.
Figure 6: A 5-node GRN example.
Figure 7: A demonstration of approximative fixpoint.
Appendix C Framework Overview
@@ 6 $\delta_{\mathcal{KB}}\leftarrow{\bm{X}}_{u}({\bm{R}}_{+}^{(k)}-{\bm{R}}_{-}^{(k)})$;
8 $(f_{y},f_{a})\leftarrow\mathrm{train}({\bm{X}}_{u},\bar{{\bm{Y}}}_{u})$;
@@ 9 $\mathcal{KB}\leftarrow\mathrm{refine}({\bm{X}}_{u},\bar{{\bm{Y}}}_{u},\mathcal{KB})$
Appendix D Experiment Details
D.1 Scalability
All experiments Figure 3 were conducted on a Slurm-managed Linux cluster equipped with Intel Xeon Gold 6342 CPUs (2.80 GHz, 32 GB system memory) and NVIDIA A100 GPUs (80 GB GPU memory). Training ALIGNED took an average of 10-12 hours per run. This demonstrates the practical scalability of ALIGNED on genome-scale problems with comprehensive knowledge bases.
D.2 Reproducibility and Hyper-parameters
@@
In Section 4.1, key hyper-parameters of our ALIGNED method as follows:

### v2

@@ 4.2 Knowledge Refinement of Gene Regulatory Networks (Q2)
We used the accuracy of interactions to test if ALIGNED’s knowledge refinement can re-discover underlying regulations from synthetic data generated from OmniPath GRN (Türei et al., 2016). For topology, we evaluated the method’s ability to produce well-structured GRNs, in terms of: (i) network modularity, for clustering quality of functional modules (Alon, 2007) and (ii) degree assortativity, for regulatory hub structures (Segal et al., 2003). In addition, we examined the method’s ability to re-discover biologically meaningful interactions by cross-referencing external pathway databases. We compared overlaps with refined pathways using a gene set recovery algorithm (Huang et al., 2018) to obtain pathway enrichment scores.
@@
The non-sparse regularized baseline ablated the refinement mechanism in Equation 6, replacing the $l_{1}$ sparse regularization term with Frobenius regularization. Existing approaches, such as GRN inference, treat the knowledge base as static instead of performing incremental refinement, and therefore are not suitable for the comparison.
@@
Observation 3. In Figure 4c, there are no significant differences of enrichment scores between the original and refined GRNs in most pathways. This indicates that ALIGNED can re-discover biologically meaningful knowledge annotated in cross-reference databases.
@@ 4.3 Perturbation Prediction on Bacterial Genome (Q3)
We evaluate our method on the Escherichia coli (E. coli) K-12 MG1655 strain using a combined dataset that includes 70 knockout perturbations by Lamoureux et al. (2023) (comprising 4 triple, 7 double, and 59 single perturbations, totaling 433 samples); and 7 data series with 16 single overexpression perturbations (73 samples) from the NCBI sequence read archive (Sayers et al., 2025). The knowledge base is constructed from the EcoCyc GRN (Moore et al., 2024), covering 315 regulator genes and 3,004 regulated genes. To evaluate generalization on unseen instances, we split the test set of unseen perturbations including 4 single overexpressions, 5 double knockouts, and 2 triple knockouts.
We tracked performance through a progressively richer procedure of ALIGNED iterations, including the neural-only models trained solely on labelled data and two complete iterations of neuro-symbolic alignment and knowledge refinement.
@@
Table 1: Performance of ALIGNED on E.coli genome. The symbol “A” indicates an alignment stage, and “R” represents a refinement stage. ALIGNED with “A-R” executes alignment and then refinement, whereas “A-R-A” performs an additional alignment step.
Model ALIGNED Stages Data Cons. Knowledge Cons. Balanced Cons.
MLP MLP only $0.3312_{\pm 0.0004}$ $0.3371_{\pm 0.0009}$ $0.3341_{\pm 0.0006}$
GNN GNN only $0.3773_{\pm 0.0026}$ $0.3605_{\pm 0.0024}$ $0.3689_{\pm 0.0012}$
ALIGNED (MLP) A $0.3872_{\pm 0.0009}$ $0.3708_{\pm 0.0006}$ $0.3784_{\pm 0.0002}$
A-R $0.3872_{\pm 0.0009}$ $0.3836_{\pm 0.0067}$ $0.3851_{\pm 0.0033}$
A-R-A $0.3812_{\pm 0.0029}$ $0.4025_{\pm 0.0020}$ $0.39404_{\pm 0.0018}$
A-R-A-R $0.3812_{\pm 0.0029}$ $0.4045_{\pm 0.0038}$ $0.3912_{\pm 0.0013}$
ALIGNED (GNN) A $0.3876_{\pm 0.0060}$ $0.3714_{\pm 0.0269}$ $0.3800_{\pm 0.0159}$
A-R $0.3876_{\pm 0.0060}$ $0.4520_{\pm 0.0070}$ $0.4130_{\pm 0.0059}$
A-R-A $0.3878_{\pm 0.0064}$ $0.4668_{\pm 0.0235}$ $0.4124_{\pm 0.0042}$
A-R-A-R $\mathbf{0.3878}_{\pm 0.0064}$ $\mathbf{0.5348}_{\pm 0.0069}$ $\mathbf{0.4288}_{\pm 0.0063}$
4.4 Ablations
In Figure 5, we compare three progressively richer variants: (i) neural-only model with no symbolic refinement (GNN only) (ii) adaptive neuro-symbolic alignment without refinement (“A”) and (iii) the complete ALIGNED framework (“A-R”). Setting (i) is a GNN trained solely on labelled data, with no symbolic alignment or refinement. Setting (ii) adds the alignment procedure on top of the neural-only baseline, but without updating the knowledge base. Setting (iii) includes both adaptive alignment and knowledge refinement steps. These ablations isolate the contribution of each component. We observe that adaptive alignment alone improves knowledge consistency, showing that the adaptor effectively incorporates information from the symbolic component. In addition, symbolic refinement provides a substantial additional improvement in knowledge consistency, while maintaining accuracy comparable to the neural-only model.
Table 2 reports ablations on $\mathcal{KB}$ and the adaptor loss (Equation 4) using the Norman et al. dataset and a GNN as the neural component. We compare ALIGNED with: (i) a randomly structured GRN as $\mathcal{KB}$; (ii) removal of the inconsistency term $\mathrm{Inc}({\bm{a}},{\bm{x}},\hat{{\bm{y}}},\mathcal{KB})$; (iii) removal of the knowledge-usage restriction term $L_{len}({\bm{a}})$; and (iv) a random weight vector $\bm{w}$}. The results show: (i) without a meaningful prior knowledge, the adaptor defaults to neural predictions; (ii) the inconsistency term encourages stronger alignment with knowledge, at a small cost in predictive accuracy; (iii) the usage-restriction term prevents over-reliance on symbolic predictions; and (iv) the weight vector helps balance predictive performance with appropriate knowledge usage.
[…]
Figure 5: Performance of the complete ALIGNED framework built with GNN. The symbol “A” indicates an alignment stage, and “R” represents a refinement stage. ALIGNED with “A-R” executes alignment and then refinement, whereas “A-R-A” performs an additional alignment step.
Table 2: Ablation of $\mathcal{KB}$ and $L_{{\bm{a}}}$. “A” indicates ALIGNED performs one alignment step.
Ablation Settings ALIGNED Stages Data Cons. Knowledge Cons. Balanced Cons.
Default A $0.5101$ $0.4172$ $0.4567$
Random $\mathcal{KB}$ A $0.5611$ $0.3190$ $0.3921$
Remove $\mathrm{Inc}$ A $0.5399$ $0.3469$ $0.4127$
Remove $L_{len}$ A $0.5047$ $0.4442$ $0.4715$
Random $\bm{w}$} A $0.4074$ $0.6194$ $0.4813$
@@ 5 Related Work
Perturbation Response Prediction. Recent approaches fall into two categories: methods that utilize the compositional nature of genetic perturbations in learning latent representations (Lotfollahi et al., 2023; Cui et al., 2024; Hao et al., 2024), hybrid methods that leverage prior knowledge from biological networks (Roohani et al., 2024; Wenkel et al., 2025; Littman et al., 2025) or textual embeddings (Wang et al., 2024). In contrast, ALIGNED does not assume GRNs to be static and can systematically refine GRNs by adaptive learning from datasets and knowledge bases to leverage reliable information.
Neuro-Symbolic Learning. The Abductive Learning (ABL) framework (Zhou, 2019) integrates deep learning with symbolic constraints through consistency optimization. Extensions include $Meta_{abd}$ (Dai & Muggleton, 2021) for visual-symbolic reasoning, $ABL_{NC}$ for knowledge refinement (Huang et al., 2023), and $ABL_{refl}$ (Hu et al., 2025) for efficient neuro-symbolic integration using reinforcement learning mechanisms. Additionally, Cornelio et al. (2023) proposed a learnable trade-off mechanism between data and knowledge sources. In comparison, ALIGNED does not assume the knowledge base to be fixed or absolutely correct. Since both biological data and curated knowledge exhibit domain-specific noise and incompleteness, ALIGNED jointly aligns neural and symbolic predictions to leverage the most reliable signals and actively refines knowledge bases during training.
@@ 6 Conclusion and Future Work
While we acknowledge the limitations in our regulatory network modelling, alternative methods (Covert et al., 2004; Stoll et al., 2017; Abou-Jaoudé et al., 2016) face significant scalability issues. Future work could explore differentiable models (Faure et al., 2023) and refine them with experimental data. Furthermore, the ALIGNED framework only assumes the availability of a neural predictor and a symbolic component with consistency measurement. Therefore, ALIGNED can be extended to different biological tasks by leveraging other prior knowledge, such as protein-protein interaction networks (Rodriguez-Mier et al., 2025) and metabolic networks (Faure et al., 2023).
Acknowledgments
We thank our colleagues at the European Bioinformatics Institute (EMBL-EBI), Julio Saez-Rodriguez, Aurelien Dugourd, Ricardo O. Ramirez-Flores, Pablo Rodríguez Mier, Philipp Schaefer, and Daniele Bottazzi, for their valuable feedback and discussions throughout the development of this work; our colleagues at Nanjing University, Yu-Feng Li, Xin-Hao Zhu, Jun-Tong Wang, and Jing He, for their discussions and suggestions throughout the initialization and implementation of this work; our collaborators at Imperial College London, Geoff S. Baldwin, Shi-Shun Liang, and Alfie J. Brown on discussions and directions on biological applications.
Reproducibility
Details of experiments and reproducibility are included in Section E. The code repository is available at https://github.com/yfxiang0112/Aligned.
@@ Appendix A Usage of Large Language Models
We used Claude and ChatGPT mainly to polish the language after all intellectual content has been drafted, along with Grammarly as a language editing tool.
@@ B.2 Demonstration
Figure 6: A 5-node GRN example.
Figure 7: A demonstration of approximative fixpoint.
@@ B.3 Discussion
Our modelling of regulatory effect given by Equation B.1 does not assume that activating and inhibitory effects combine through simple additive rules. The formulation is path-dependent and encodes how positive and negative regulations compose across multi-step pathways. However, such modelling is still not enough to describe the biological reality, and future work could further explore the other differentiable modelling approaches of genome-scale GRN under the ALIGNED framework.
In actual experiments in Section 4.1, we introduced an assumption that up/down regulation of a node can be decided by its in-degree of positive and negative interactions, in order to capture more detailed regulatory behaviours. This is achieved by using an integer variant of Equation 2.2. In this setting, the value of node 5 in Figure 7 will be -1 when $k=2$, i.e. $\delta_{\langle{\bm{R}}_{+}^{(2)},{\bm{R}}_{-}^{(2)}\rangle}({\bm{x}})=[0,-1,1,1,-1]^{\top}$, and the network behaviour will be more complicated in complex networks.
Appendix C Discussions of Convergence
The convergence of the ALIGNED iterations is guaranteed under certain theoretical assumptions. The inconsistency measurement defined in Equation 2 is non-negative, lower-bounded by 0. Since both the neural and the symbolic component are updated with the neuro-symbolic prediction $\langle{\bm{X}}_{u},\bar{{\bm{Y}}_{u}}\rangle$ in each iteration, the inconsistency measurement reduces if the following assumptions hold: (a) the neural prediction $\hat{{\bm{Y}}_{u}}$ aligns better with $\bar{{\bm{Y}}_{u}}$ after update, and (b) the symbolic prediction $\delta_{\mathcal{KB}}({\bm{X}}_{u})$ aligns better with $\bar{{\bm{Y}}_{u}}$ after the knowledge refinement. Therefore, the inconsistency decreases monotonically until it reaches a fixed point, where further updates to either module no longer reduce disagreement and the ALIGNED reaches convergence.
In practice, ALIGNED also converges when the assumptions (a), (b) do not strictly hold in theory. Empirically, the iterations are observed to stabilize within a small number of iterations (typically 2-3), as shown in Figure 5.
Appendix D Framework Overview
@@ 6 $\delta_{\mathcal{KB}}\leftarrow{\bm{X}}_{u}({\bm{R}}_{+}^{(k)}-{\bm{R}}_{-}^{(k)})$;
8 $(f_{y},f_{a})\leftarrow\mathrm{update}({\bm{X}}_{u},\bar{{\bm{Y}}}_{u})$;
@@ 9 $\mathcal{KB}\leftarrow\mathrm{refine}({\bm{X}}_{u},\bar{{\bm{Y}}}_{u},\mathcal{KB})$
Appendix E Experiment Details and Reproducibility
E.1 Reproducibility and Code Availability
@@
E.2 Hardware and Runtime
All experiments were conducted using a single Intel Xeon Gold 6342 CPU (2.80 GHz, 32 GB memory) and a single NVIDIA A100 GPU (80 GB). Training ALIGNED took an average of 10-12 hours per run. Under our experimental setup, the training time of ALIGNED is manageable given the size of the GRN and the complexity of the refinement procedure.
E.3 Hyperparameters
In Section 4.1, key hyperparameters of our ALIGNED method as follows:
@@ E.3 Hyperparameters
The weight vector $\bm{w}$} is computed separately for each dataset, but always using the same fixed procedure above. In particular, the component of $\bm{w}$} that reflects data-knowledge inconsistency is derived only from the training split, while the other components encode priors from the structured knowledge bases. This makes $\bm{w}$} a dataset-specific but non-hand-tuned quantity.
@@
E.4 Experiments on Hyperparameters of the Adaptor
To study the effect on ALIGNED’s overall performance of each hyperparameter, we conducted a comparison of different parameter selection of $k$ in Equation 2.2, $\theta$ in Equation 4, $t_{0},t_{k}$ in Equation 6, as well as different computing for $\bm{w}$} in Equation 4. The comparison in Table LABEL:tab:hyperparam was ran for 2 ALIGNED loops, and we report the performance at neuro-symbolic alignment of the first loop (Align 1) since the hyperparameters put most effect on the behavior of the adaptor, as well as the performance at the end of the loops (Refine 2).
All comparisons were performed on the Norman et al. (2019) dataset with GNN as the neural component, including the default setting in Section 4.1 (denoted as “default”, with $k=4$, $\theta=0.3$ and $\bm{w}$} computed as Section E.3), different values of $k,\theta$, and different coefficients in computing $\bm{w}$} (“term $i$ +” means that the $i$-th term’s coefficient in Section E.3 $\bm{w}$}’s formula is increased by $0.1$, following by clipping the $\bm{w}$}’s elements to keep in $[0,1]$).
The results in Table LABEL:tab:hyperparam indicate that, maximum allowed pathway length ($k$) in the modelling of $\mathcal{KB}$ is positively correlated with data consistency (predictive performance on test data), but negatively correlated with the knowledge consistency. The threshold of $\mathcal{KB}$ usage ($\theta$) has a minor effect on encouraging $\mathcal{KB}$ usage with higher $\theta$. Term coefficients in $\bm{w}$}’s formula has a positive effect on balanced consistency, in comparison with setting all elements of $\bm{w}$} to 0.5. Every term in $\bm{w}$}’s formula is positively correlated with knowledge consistency but slightly compromises data consistency.
Table 3: Performance of ALIGNED on different hyperparameter selections. The symbol “A” indicates an alignment stage, and “R” represents a refinement stage. ALIGNED with “A” executes one alignment step, whereas “A-R-A-R” performs two rounds of alignment and refinement steps.
ALIGNED Stages Operation Data Cons. Knowledge Cons. Balanced Cons.
GNN only default $0.5344$ $0.3823$ $0.4397$
A default $0.5101$ $0.4172$ $0.4567$
$k=0$ $0.4426$ $0.3270$ $0.3719$
$k=2$ $0.4960$ $0.4545$ $0.4738$
$k=6$ $0.5129$ $0.4159$ $0.4568$
$\theta=0.1$ $0.5148$ $0.4051$ $0.4502$
$\theta=0.5$ $0.5047$ $0.4442$ $0.4715$
${\bm{w}}=\mathbf{0.5}$ $0.3944$ $0.6381$ $0.4744$
$\bm{w}$}: term 1 + $0.4993$ $0.6577$ $0.5624$
$\bm{w}$}: term 2 + $0.5120$ $0.4335$ $0.4678$
$\bm{w}$}: term 3 + $0.4999$ $0.4294$ $0.4606$
A-R-A-R default $0.5233$ $0.6570$ $0.5788$
$k=0$ $0.4377$ $0.3388$ $0.3788$
$k=2$ $0.5115$ $0.6416$ $0.5656$
$k=6$ $0.5028$ $0.6606$ $0.5658$
$\theta=0.1$ $0.5254$ $0.6549$ $0.5795$
$\theta=0.5$ $0.5175$ $0.6607$ $0.5761$
${\bm{w}}=\mathbf{0.5}$ $0.4047$ $0.6476$ $0.4853$
$\bm{w}$}: term 1 + $0.5003$ $0.6509$ $0.5609$
$\bm{w}$}: term 2 + $0.5226$ $0.6614$ $0.5798$
$\bm{w}$}: term 3 + $0.5099$ $0.6582$ $0.5700$
E.5 Experiments on Hyperparameters of Knowledge Refinement
We conducted experiment in Section 4.2 to examine the effect of the approximation coefficient $t_{k}$ and $t_{0}$ in Equation 6. The default setting in Section 4.2 ($t_{0}=100$$t_{k}=1$) is denoted as “Default”, and the performance was compared in $F_{1}$ score of direct interactions, same with Figure 4a. As shown in Table LABEL:tab:refine_t, the performance of knowledge refinement is positively correlated with $t_{0}$ and $t_{k}$. Larger values of $t_{0}$ and $t_{k}$ than default have diminishing improvement returns in performance. Generally, larger values of $t_{0}$ and $t_{k}$ require higher computational resource. Therefore, we decided on the current default values based on both performance and computational cost.
Table 4: Performance of knowledge refinement on different hyperparameter selections
Noisy Interactions Default $t_{0}=1000$, $t_{k}=10$ $t_{0}=10$, $t_{k}=0.1$
0% $0.8896$ $0.9052$ $0.6726$
5% $0.8554$ $0.8895$ $0.6262$
10% $0.8339$ $0.8753$ $0.5921$
30% $0.7894$ $0.8181$ $0.5347$
50% $0.7233$ $0.7604$ $0.5357$
E.6 Dataset Split
The detail of train-test data split in Experiment of Section 4.1 is shown in Table LABEL:tab:split.
Table 5: Test dataset split in Experiment of Section 4.1
Dataset Replicates Test set perturbations
Norman - FEV, TBX3, AHR, S1PR2, TBX2, MAP7D1, ZBTB1, OSR2, ISL2, BPGM, SGK1, MAP7D1, PRDM1, UBASH3B + OSR2, PRDM1 + CBFA2T3, SGK1 + TBX3, TBX3 + TBX2, PTPN12 + OSR2, BPGM + SAMD1, AHR + KLF1, FEV + MAP7D1, SAMD1 + ZBTB1, CEBPB + OSR2, SGK1 + S1PR2, FEV + ISL2, SGK1 + TBX2, ETS2 + MAP7D1, AHR + FEV, FEV + CBFA2T3, FOSB + OSR2, BPGM + ZBTB1
Dixit 1 ELF1, E2F4, CIT, CEP55, AURKA
2 ELK1, NR2C2, ECT2, CENPE
3 ELK1, E2F4
4 CENPE, TOR1AIP1, RACGAP1
5 GABPA, E2F4, CENPE
Adamson 1 DDIT3, OST4, SEC61A1, HARS, IARS2, SOCS1, MRGBP, MTHFD1, TMEM167A, TTI2, SLC39A7, MRPL39, SEC63, MARS, ARHGAP22, EIF2S1, TIMM44, PTDSS1, CHERP, SRPRB, DHDDS, COPZ1, CARS, P4HB
2 BHLHE40, DDIT3, SAMM50, MTHFD1, EIF2B2, PSMD4, EIF2S1, CCND3, PPWD1, COPZ1, EIF2B3, XRN1, TMED2, SPCS2
3 DASAMM50, MRGBP, TMEM167A, NEDD8, HSPA9, MRPL39, FARSB, HSD17B12, SEC61B, ASCC3, ARHGAP22, HSPA5, CHERP
4 ZNF326, DDIT3, OST4, QARS, MRGBP, MRPL39, SEC63, MANF, HSD17B12, SARS, TIMM44, DDOST, SLC35B1, CCND3, EIF2B3, XRN1, TMED2, SPCS3
5 ZNF326, BHLHE40, SEC61A1, IER3IP1, YIPF5, IARS2, ATP5B, MTHFD1, GNPNAT1, TARS, PSMD4, CHERP, DHDDS, DNAJC19
Appendix F Additional Results
F.1 Curation Effort and $\mathcal{KB}$ Prediction Contribution
We conducted an experiment to stratify edges in OmniPath by their curation effort (a proxy for knowledge reliability) and measured how often the adaptor selects the symbolic versus neural predictions in each regime. ALIGNED did not have access to this curation effort information during training. Our results in Table LABEL:tab:curation_effort show that the adaptor automatically down-weights the symbolic predictions when the GRN is less reliable (low curation effort) and relies more heavily on the neural predictor in those cases. Conversely, when the edges in $\mathcal{KB}$ are better curated, the adaptor increases its symbolic usage.
Table 6: Knowledge vs. data usage across curation effort levels
Curation Effort Knowledge Usage Data Usage
0 13.3% 86.7%
[1,5) 24.7% 75.3%
[5,10) 31.9% 68.1%
$\geq 10$ 31.7% 68.3%
F.2 GRN failure modes
We experimented using multiple external knowledge sources built by Chevalley et al. (2025), include STRINGdb (PPI), CORUM (PPI) and a ChipSeq network derived from Chip-Atlas and ENCODE, as independent cross-references.
We stratified gene pairs into three categories: (i) not regulated in both (sparse), (ii) regulated in both (high confidence), (iii) regulated only in our $\mathcal{KB}$ or in reference (low confidence). The rationale is that interactions supported by both our $\mathcal{KB}$ (OmniPath) and an external reference are more likely to be reliable, whereas interactions found in only one source may imply biases, incompleteness, or systematic errors. The results in Table 7 demonstrate that when coverage is sparse or inconsistent, ALIGNED relies heavily on the neural component ( 90%). This prevents the model from propagating uncertain symbolic signals. In addition, when prior knowledge is less reliable, the symbolic component contributes less to the overall predictions compared to high-confidence edges.
Table 7: Knowledge vs. data usage comparing with external databases
Database Edges Knowledge Usage Data Usage
STRINGdb sparse 10.5% 89.5%
high confidence 74.6% 25.4%
low confidence 54.8% 45.2%
CORUM sparse 89.4% 10.6%
high confidence 16.9% 83.1%
low confidence 28.2% 71.8%
ChipSeq sparse 89.4% 10.6%
high confidence 25.4% 74.6%
low confidence 29.4% 70.6%
F.3 Interaction Recovery
We randomly sampled interactions (3 replicates) from our $\mathcal{KB}$ that simultaneously exists in a PPI database (STRING or Corum) and the ChipSeq network by Chevalley et al. (2025), which we believe that they tend to be true interactions with a high confidence. We removed them selected interactions from our $\mathcal{KB}$, ran ALIGNED using the $\mathcal{KB}$ with removed edges and examined whether the removed interactions can be recovered after 2 alignment / refinement iterations.
The selected interactions and the performance of recovery are listed in Table 8. The results show that the ALIGNED iterations are able to correct potential noise and incompleteness in $\mathcal{KB}$ using neural predictions.
Table 8: Recovery rate of confident interaction removal
Replicate Recovered/Total Interactions
1 11/12 JUN - CTNNB1, SPI1 - FCGRT, CEBPA - SPI1, CEBPB - TGFB1, CEBPB - GDF15, CEBPB - TRIB3, CEBPB - HSP90AA1, CEBPB - RUN X2, CEBPA - CEBPE, JUN - BEX2, JUN - GSTP1, JUN - KRT18
2 10/10 JUN - CTNNB1, JUN - GSTP1, CEBPB - TGFB1, CEBPB - CEBPA, CEBPB - TRIB3, CEBPB - HSP90AA1, CEBPB - DDIT3, CEBPB - GST P1, CEBPB - CSF3R, CEBPA - CEBPE
3 10/11 JUN - STMN1, JUN - MYC, JUN - GSTP1, JUN - VIM, JUN - KRT18, SPI1 - FCGRT, CEBPB - CSF3R, CEBPB - RUNX2, CEBPB - TRIB3, CEBPB - GDF15, CEBPA - CSF3R
F.4 Cross-Reference for Section 4.2
We extended the GRN refinement experiment in Section 4.2 with additional cross-reference comparisons using external databases built by Chevalley et al. (2025). Performance of the knowledge refinement is measured via comparing the accuracy of GRN after refinement contrasting with the reference database, in terms of $F_{1}$ score. The performance is compared with the non-sparse baseline (using Frobenius regularization), as noted in Section 4.2. Our results are shown in Table 9, The results demonstrate that our method is able to learn interactions that are verified by external reference sources.
Table 9: GRN after refinement contrasting with reference databases in terms of $F_{1}$
Method Noisy Interactions STRINGdb CORUM ChipSeq
ALIGNED $0\%$ 0.4561 0.0310 0.2015
$5\%$ 0.4406 0.0293 0.1814
$10\%$ 0.4271 0.0270 0.1764
$30\%$ 0.4093 0.0258 0.1714
$50\%$ 0.3706 0.0207 0.1363
Non-sparse $0\%$ 0.2644 0.0058 0.0372
$5\%$ 0.2600 0.0054 0.0350
$10\%$ 0.2685 0.0074 0.0530
$30\%$ 0.2351 0.0041 0.0315
$50\%$ 0.0472 0.0002 0.0018
F.5 Shortcut Regulations in Section 4.2
To examine whether the knowledge refinement learns undesirable shortcut signals from the data, we extended the GRN-refinement experiment in Section 4.2 by comparing the proportion of direct versus indirect interactions after knowledge refinement on the noisy GRN. We evaluated how many direct and indirect interactions in the original GRN were learned, where direct interactions are desired and indirect interactions represents shortcut regulatory effects in the data.
The results in Table 10 show that ALIGNED’s refinement predominantly recovers direct interactions, while indirect ones remain limited. In contrast, the non-sparse baseline (using Frobenius regularization) overfits to indirect interactions. These findings suggest that our sparse refinement strategy effectively guards against overfitting to shortcut signals.
Table 10: Direct vs. indirect interactions in Section 4.2
Noisy Interactions ALIGNED Non-sparse
Direct Indirect Direct Indirect
0% 81.4% 17.0% 16.6% 68.0%
5% 73.3% 21.3% 15.0% 68.6%
10% 71.8% 23.2% 24.1% 53.6%
30% 71.2% 18.8% 11.9% 70.5%
50% 58.6% 24.7% 0.6% 36.8%